# The effect of self-induced Marangoni flow on polar-nematic waves in active-matter systems

### v1

@@ 1 Introduction
To model the characteristics of “dry” systems, where the motion of the fluid matrix supporting the self-propelled agents is not considered, the density distribution of the agents is described by the Boltzmann equation. It accounts for the microscopic rules governing binary collisions and random reorientations of the agents [12, 13, 14]. In the small-velocity regime, truncated forms of the Boltzmann equation can be derived by assuming dominant polar [12] or polar and nematic order parameter fields [13, 14]. Analyzing the resulting set of closed hydrodynamic equations for these fields reveals that the observed spatio-temporal patterns in active matter originate in spontaneous symmetry breaking that occurs when the agent density surpasses a critical threshold [14].
The situation and resulting models are more complex for “wet” systems, defined by a two-way coupling of the motion of the self-propelled agents and the surrounding fluid medium. In this scenario, the Boltzmann equation for the agents’ density distribution is coupled to the Navier-Stokes equations that govern fluid motion. Here, we concentrate on a two-dimensional (2D) colony of self-propelled agents bound to move along the planar surface of a thin viscous film of liquid on a solid substrate. Then, the Boltzmann equation is extended by an advection term representing liquid flow at the film surface. Previously, we investigated the dynamics of a colony of noninteracting self-propelled agents on the surface of a liquid film with a deformable interface, particularly focusing on self-induced Marangoni flow [15, 16]. Such flows result from surface tension gradients that are due to concentration gradienst of the agents.
When the average orientation of the self-propelled agents has a component orthogonal to the free surface, their motile force exerts an excess pressure similar to the hydrostatic pressure for a pending liquid film [15]. When the excess pressure dominates the stabilizing surface tension, a Rayleigh-Taylor instability occurs [17, 18]. It is a long-wave stationary instability, i.e. its onset occurs at zero wavenumber and perturbations grow monotonically. In the classification of Ref. [19] this corresponds to a Cahn-Hilliard type. In a passive system (i.e. one that approaches thermodynamic equilibrium) such an instability results in the formation of drops that slowly coarsen. However, the situation changes if also a component of motile force exist that is parallel to the free surface. This results in the generation of intricate self-propelled dynamic density patterns and is studied in Ref. [16] by means of coupled thin-film and Smoluchowski equations. Note that in the context of the dynamics of an active Brownian particle, the Smoluchowski equation is essentially identical to the Boltzmann equation with an additional condition for the translational diffusivity constant which is linked to the temperature via the fluctuation-dissipation theorem [20].
@@
Here, we consider a model system of swimmers moving parallel to the undeformable free surface of a liquid film on a solid substrate, and study the dynamics of the relevant order parameter fields. In this way we explore how the advective flow of the liquid matrix, induced by surface tension gradients affects the propagation of density waves. We expand a macroscopic continuum approach developed for dry active systems with local nematic alignment [13, 14]. In particular, we consider the wet active system that results from the incorporation of self-induced Marangoni flow into the hydrodynamic mean-field equations for the dominant polar and nematic order parameter fields.
Specifically, we demonstrate that upon gradually increasing the Marangoni parameter, which characterises the relative strength of the Marangoni flow as compared to the self-propulsion speed of the agents, density waves broaden while their speed may either increase or decrease depending on wavelength and on overall mean density. This result echos the influence of self-induced Marangoni flow on travelling reaction-diffusion waves confined in a liquid drop or film [24, 25]. There, the reaction is fueled by an exothermic autocatalytic chemical reaction, such as the Belousov-Zhabotinsky or the iodate-arsenous acid reaction, and all reactants are non-motile. Similar to our observations, the inclusion of the Marangoni flow in autocatalytic chemical reaction systems leads to a change in propagation speed and front deformation. We also demonstrate two different bifurcation scenarios that result in the destruction of the waves at a relatively strong Marangoni flow. In the first scenario, a gradual increase in the Marangoni parameter results in the approach and ensuing collision of branches of stable und unstable finite-amplitude waves. The corresponding saddle-node bifurcation results in a discontinuous disappearance of the density waves. In the second scenario, a gradual increase in the the Marangoni parameter results in a gradual decrease of the wave amplitude until it eventually vanishes in a continuous transition via a wave bifurcation, i.e. a finite-wavelength Hopf bifurcation, of a uniform state.
@@ 2 Governing equations
Consider a colony of self-propelled agents that are bound to move along a planar surface of a thin liquid carrier film of uniform and constant thickness $h$ and dynamic viscosity $\mu$ that is supported by a solid bottom plate. The orientation of all agents remains parallel to the film surface at all times. Any two agents interact with each other via local binary collision rules. Thereby, any two colliding agents will align along their average pre-collision orientation when colliding at an acute angle and anti-align when colliding at an obtuse angle. Here, we adopt the kinetic Boltzmann approach developed for dry active matter in the absence of a fluid matrix [13, 14]. The evolution of the one-particle probability density distribution $\rho({\bm{r}},\theta,t)$ that depends on the position in the substrate plane ${\bm{r}}=(x,y)$ and the orientation $\theta$ of the agents is described by the kinetic Boltzmann equation
$\partial_{t}\rho+{\bm{\nabla}}\cdot[({\bm{U}}+v_{0}{\bm{e}}(\theta))\rho]=d{\bm{\nabla}}^{2}\rho+I_{\text{rot}}+I_{\text{coll}},$ (1)
[…]
@@
Here, we assume that the deformation of the film surface can be neglected, which is valid at small capillary numbers $\mu U/\sigma_{0}\ll 1$, where $\sigma_{0}$ is the surface tension of the bare liquid in the absence of surfactant [27]. However, at large capillary number a complete description of the dynamics of a wet system would necessarily include the dynamics of film surface deformations. This will be discussed in a future study.
The velocity ${\bm{U}}$ of the Marangoni flow at the planar film surface depends nonlocally on gradients of the surfactant density $c$. It can be represented in the form [27]
${\bm{U}}=-\frac{Eh}{4\pi^{2}c_{0}\mu}\int d^{2}k\frac{i{\bm{k}}\hat{c}(\sinh^{2}(|{\bm{k}}|h)-(|{\bm{k}}|h)^{2})}{|{\bm{k}}|h\sinh(2|{\bm{k}}|h)-2(|{\bm{k}}|h)^{2}}e^{i{\bm{k}}\cdot{\bm{r}}},$ (2)
@@
where $E=-c_{0}(d\sigma(c)/dc)\big{|}_{c=c_{0}}>0$ is the Marangoni elasticity of the film surface that encodes the strength of change of the surface tension $\sigma(c)$ with $c$. Further, ${\bm{k}}=(k_{x},k_{y})$ is a 2D wave number and $\hat{c}({\bm{k}},t)=4\pi^{2}\int d^{2}r\,c({\bm{r}},t)e^{-i{\bm{k}}\cdot{\bm{r}}}$ is the Fourier transform of $c$.
@@
In long-wave approximation [17, 28, 29], valid when the thickness of the layer is much smaller than the characteristic horizontal length scale of the variation in surfactant density, i.e. $h\ll 2\pi/|{\bm{k}}|$, Eq. (2) reduces to the well-known local relationship [17]
${\bm{U}}({\bm{r}},t)=-\frac{Eh}{4c_{0}\mu}{\bm{\nabla}}c({\bm{r}},t).$ (3)
The uniform disordered state $\rho({\bm{r}},\theta)=1/(2\pi)$ may become unstable to spatial and orientational perturbations, leading to a transition toward dynamic inhomogeneous ordered states including steady clusters or bands, and propagating waves.
To approximate the Boltzmann equation (1), next we introduce order parameter fields
[…]
and assume the scaling and truncation
$\partial_{t}\sim\epsilon,~{}~{}\partial_{x,y}\sim\epsilon,~{}~{}c-c_{0}\sim\epsilon,$
$\rho_{1,2}\sim\epsilon,~{}~{}\rho_{3,4}\sim\epsilon^{2},~{}~{}\rho_{q>4}=0$ (5)
@@
where $\epsilon\ll 1$ is the dimensionless super-criticality, i.e. the distance from the onset of the order-disorder transition. We follow Ref. [13, 14] and derive a hydrodynamic approximation consisting of kinetic equations for the polar ($\rho_{1}$) and nematic ($\rho_{2}$) order parameter fields. The scaling (5) automatically restricts the applicability of the hydrodynamic approximation to relatively small magnitudes and slow variations of $\rho_{1}$ and $\rho_{2}$. In consequence, the approximation is valid only in the vicinity of the transition point where the disordered state $c=c_{0}$ becomes unstable.
Noticing that the Marangoni velocity ${\bm{U}}({\bm{r}},t)$ is independent of the orientation of the agents and further assuming that it is of the same order of magnitude as the self-propulsion speed $v_{0}$, the hydrodynamic approximation derived in Refs. [13, 14] can be directly extended to include the Marangoni flow. The resulting closed system of coupled advection-diffusion-reaction equations for the real-valued density $c$ and the complex-valued polar and nematic order parameter fields $\rho_{1}$ and $\rho_{2}$ is given by
$\partial_{t}c+{\bm{\nabla}}\cdot({\bm{U}}c)$ $=$ $-\frac{v_{0}}{2}\left(\partial\rho_{1}^{*}+\partial^{*}\rho_{1}\right)+d{\bm{\nabla}}^{2}c$
$\partial_{t}\rho_{1}+{\bm{\nabla}}\cdot({\bm{U}}\rho_{1})$ $=$ $-(\alpha_{0}+c\alpha_{1})\rho_{1}+\alpha_{2}\rho_{1}^{*}\rho_{2}-\alpha_{3}|\rho_{2}|^{2}\rho_{1}$ (6)
-- $\frac{v_{0}}{2}(\partial c+\partial^{*}\rho_{2})+\gamma_{1}\rho_{2}^{*}\partial\rho_{2}+d{\bm{\nabla}}^{2}\rho_{1}$
$\partial_{t}\rho_{2}+{\bm{\nabla}}\cdot({\bm{U}}\rho_{2})$ $=$ $(-\beta_{0}+c\beta_{1})\rho_{2}+\beta_{2}\rho_{1}^{2}-\beta_{3}|\rho_{2}|^{2}\rho_{2}$
@@
-- $\gamma_{3}\rho_{1}^{*}\partial\rho_{2}-\gamma_{4}\partial^{*}(\rho_{1}\rho_{2})+d{\bm{\nabla}}^{2}\rho_{2}.$
Here, $\partial=\partial_{x}+i\partial_{y}$ and ${\bm{\nabla}}=(\partial_{x},\partial_{y})^{T}$ while the star indicates the complex conjugate. The system (6) is closed by Eq. (2) [or Eq. (3) in the long-wave limit] that relates ${\bm{U}}$ to the number density $c$. Note that we have retained translational diffusion with diffusivity $d$ to ensure numerical stability and consistency with the Marangoni term. For ease of comparison, system (6) is nondimensionalised using the scaling introduced in Ref. [14] for a dry system, i.e. when ${\bm{U}}=0$. Namely, time is scaled with the characteristic duration of the ballistic flight path of the agents $\tau=r^{-1}$, where $r$ is the rate of random orientation changes of the agents. The spatial coordinates are scaled with $v_{0}\tau$, and the one-particle density distribution $\rho$ is measured in units of $r/Dv_{0}$, where $D$ is the size of the agent particles. In dimensionless units, the dynamics of a wet system in the long-wave limit is described by Eqs. (6) with $v_{0}=1$ and ${\bm{U}}=-\text{Ma}{\bm{\nabla}}c$ where $\text{Ma}=Eh/(4v_{0}^{2}\mu\tau)$ represents the Marangoni parameter that controls the relative strength of the Marangoni flow as compared to the self-propulsion speed.
The coefficients $\alpha_{i}$ and $\beta_{i}$ in Eqs. (6) determine the properties of the spatially uniform states, while the $\gamma_{i}$ are responsible for the coupling between the gradients of polar and nematic order parameter fields. Following Ref. [14], we choose the coefficients $\alpha_{i}$ and $\beta_{i}$ to correspond to a system that is known to feature several transitions in absence of Marangoni flow. When the mean number density $c_{0}$ is varied at fixed values of the coefficients $\alpha_{i}$, $\beta_{i}$ and $\gamma_{i}$, transitions occur between a uniform disordered state and a uniform ordered nematic state and as well between the latter and a uniform ordered polar-nematic state of mixed symmetry. The parameter values employed in the following are
@@
$\gamma_{1}$ $=$ $3.7\displaystyle 1,~{}\gamma_{2}=0.85,~{}\gamma_{3}=0.23,~{}\gamma_{4}=$ 3.7 . (7)
The translational diffusivity is varied between $d=10^{-2}$ and $d=10^{-1}$.

### v2

@@ 1 Introduction
To model the characteristics of “dry” systems, where the motion of the fluid matrix supporting the self-propelled agents is not considered, the density distribution of the agents is described by the Boltzmann equation. It accounts for the microscopic rules governing binary collisions and random reorientations of the agents [12, 13, 14]. In the small-velocity regime, truncated forms of the Boltzmann equation can be derived by assuming dominant polar [12] or polar and nematic order parameter fields [13, 14]. Analysing the resulting set of closed hydrodynamic equations for these fields reveals that the observed spatio-temporal patterns in active matter originate in spontaneous symmetry breaking that occurs when the agent density surpasses a critical threshold [14].
The situation and resulting models are more complex for “wet” systems, defined by a two-way coupling of the motion of the self-propelled agents and the surrounding fluid medium. In this scenario, the Boltzmann equation for the agents’ density distribution is coupled to the Navier-Stokes equations that govern fluid motion. Here, we concentrate on a two-dimensional (2D) colony of self-propelled agents bound to move along the planar surface of a thin viscous film of liquid on a solid substrate. Then, the Boltzmann equation is extended by an advection term representing liquid flow at the film surface. Previously, we investigated the dynamics of a colony of noninteracting self-propelled agents on the surface of a liquid film with a deformable interface, particularly focusing on self-induced Marangoni flow [15, 16]. Such flows result from surface tension gradients that are due to concentration gradients of the agents.
When the average orientation of the self-propelled agents has a component orthogonal to the free surface, their motile force exerts an excess pressure similar to the hydrostatic pressure for a pending liquid film [15]. When the excess pressure dominates the stabilising surface tension, a Rayleigh-Taylor instability occurs [17, 18]. It is a long-wave stationary instability, i.e. its onset occurs at zero wavenumber and perturbations grow monotonically. In the classification of Ref. [19] this corresponds to a Cahn-Hilliard type. In a passive system (i.e. one that approaches thermodynamic equilibrium) such an instability results in the formation of drops that slowly coarsen. However, the situation changes if also a component of motile force exist that is parallel to the free surface. This results in the generation of intricate self-propelled dynamic density patterns and is studied in Ref. [16] by means of coupled thin-film and Smoluchowski equations. Note that in the context of the dynamics of an active Brownian particle, the Smoluchowski equation is essentially identical to the Boltzmann equation with an additional condition for the translational diffusivity constant which is linked to the temperature via the fluctuation-dissipation theorem [20].
@@
Here, we consider a model system of swimmers moving parallel to the non-deformable free surface of a liquid film on a solid substrate, and study the dynamics of the relevant order parameter fields. In this way we explore how the advective flow of the liquid matrix, induced by surface tension gradients affects the propagation of density waves. We expand a macroscopic continuum approach developed for dry active systems with local nematic alignment [13, 14]. In particular, we consider the wet active system that results from the incorporation of self-induced Marangoni flow into the hydrodynamic mean-field equations for the dominant polar and nematic order parameter fields.
Specifically, we demonstrate that upon gradually increasing the Marangoni parameter, which characterises the relative strength of the Marangoni flow as compared to the self-propulsion speed of the agents, density waves broaden while their speed may either increase or decrease depending on wavelength and on overall mean density. This result echos the influence of self-induced Marangoni flow on travelling reaction-diffusion waves confined in a liquid drop or film [24, 25]. There, the reaction is fuelled by an exothermic autocatalytic chemical reaction, such as the Belousov-Zhabotinsky or the iodate-arsenous acid reaction, and all reactants are non-motile. Similar to our observations, the inclusion of the Marangoni flow in autocatalytic chemical reaction systems leads to a change in propagation speed and front deformation. We also demonstrate two different bifurcation scenarios that result in the destruction of the waves at a relatively strong Marangoni flow. In the first scenario, a gradual increase in the Marangoni parameter results in the approach and ensuing collision of branches of stable und unstable finite-amplitude waves. The corresponding saddle-node bifurcation results in a discontinuous disappearance of the density waves. In the second scenario, a gradual increase in the the Marangoni parameter results in a gradual decrease of the wave amplitude until it eventually vanishes in a continuous transition via a wave bifurcation, i.e. a finite-wavelength Hopf bifurcation, of a uniform state.
@@ 2 Governing equations
Figure 1: Schematic representation of the system in (a) side view and (b) top view. Swimming agents represented by filled circles move along a flat liquid-air interface of a thin liquid film of uniform thickness $h$. On a coarse-grained level, swimmers of concentration $c(x,y,t)$ act as a surfactant, inducing Marangoni flow with velocity $\mathbf{u}(x,y,z,t)$ directed away from the region with higher density. The agents are advected by the surface flow field $\mathbf{U}(x,y,t)=\mathbf{u}(x,y,z=h,t)$, shown by blue arrows in (b). The orientations of individual agents are indicated by black arrows in (b) while the coarse-grained particle density is color-coded.
@@
Consider a colony of self-propelled agents that are bound to move along a planar surface of a thin liquid carrier film of uniform and constant thickness $h$ and dynamic viscosity $\mu$ that is supported by a solid bottom plate, as schematically shown in Fig.1. The orientation of all agents remains parallel to the film surface at all times. Any two agents interact with each other via local binary collision rules. Thereby, any two colliding agents will align along their average pre-collision orientation when colliding at an acute angle and anti-align when colliding at an obtuse angle. Here, we adopt the kinetic Boltzmann approach developed for dry active matter in the absence of a fluid matrix [13, 14].
To describe the interaction with the liquid, we neglect the size of the agents and assume that the main contribution to their motion is due to the advection by the liquid along the surface of the carrier film with the local velocity $\mathbf{U}=(U_{x},U_{y})=\mathbf{u}(x,y,z=h,t)$, where $\mathbf{u}(x,y,z,t)$ is the horizontal flow field in the bulk. We emphasise that the motion of solid particles of finite size trapped at an interface between two fluids is a complex hydrodynamic problem that (in the limit of small Reynolds number) involves the solution of the Stokes equation supplemented with no-slip boundary conditions at the particle surface and the continuity conditions for the flow field involving stresses at the interfaces as well as wetting conditions. For a comprehensive review of the motion of passive colloidal particles at liquid interfaces see e.g. [26]. The complete hydrodynamic description becomes even more involved for self-propelled particles moving near penetrable liquid-liquid interfaces and can only be resolved numerically [27]. Here, we treat the particles as point-like implying that they are advected with the velocity $\mathbf{U}$.
On the continuum level, the ensemble of self-propelled agents at the film surface is described by the one-particle probability density distribution $\rho(\mathbf{r},\theta,t)$ that depends on the position in the film surface plane $\mathbf{r}=(x,y)$ (that is parallel to the solid substrate) and the orientation $\theta$ of the agents. Its time evolution is described by the kinetic Boltzmann equation
$\partial_{t}\rho+\mathbf{\nabla}\cdot[(\mathbf{U}+v_{0}\mathbf{e}(\theta))\rho]=d\mathbf{\nabla}^{2}\rho+I_{\text{rot}}+I_{\text{coll}},$ (1)
where $v_{0}$ is the self-propulsion speed, $\mathbf{e}(\theta)=(\cos(\theta),\sin(\theta))$ is a unit orientation vector, $d$ is the translational diffusivity, $I_{\text{rot}}$ describes rotational diffusion and $I_{\text{coll}}$ is the collision term that describes the alignment of the agents. If the one-particle density distribution $\rho(\mathbf{r},\theta,t)$ is normalised in such a way that $\int_{A}d\mathbf{r}\int_{0}^{2\pi}d\theta\rho(\mathbf{r},\theta,t)=N$, where $N$ is the total number of agents for a liquid film of surface area $A$, then their local number density per surface area is $c(\mathbf{r},t)=\int_{0}^{2\pi}d\theta\,\rho(\mathbf{r},\theta,t)$. Then $c_{0}=A^{-1}\int_{A}c(\mathbf{r},t)\,d\mathbf{r}=NA^{-1}$ is the mean number density.
The rotational diffusion term $I_{\text{rot}}$ in equation (1) depends on the one-particle density $\rho(\mathbf{r},\theta,t)$, while for binary collisions, the collision term $I_{\text{coll}}$ depends on the two-particle density function $\rho_{2}(\mathbf{r}_{1},\theta_{1},\mathbf{r}_{2},\theta_{2},t)$. However, in the mean field approximation one truncates the resulting hierarchy of equations assuming the closure relation $\rho_{2}(\mathbf{r}_{1},\theta_{1},\mathbf{r}_{2},\theta_{2},t)=\rho(\mathbf{r}_{1},\theta_{1},t)\rho(\mathbf{r}_{2},\theta_{2},t)$ [28]. This allows one to represent $I_{\text{coll}}$ as a quadratic functional of $\rho$, as described in Refs. [13, 14].
Equation (1) must be supplemented with an equation for the advection flow field $\mathbf{U}(\mathbf{r},t)$. Similar to Ref. [15, 16] we assume that in the presence of the carrier liquid the self-propelled agents act as a surfactant, i.e. due to the solutal Marangoni effect, gradients in their density induce a Marangoni flow at the film surface. Note that translational diffusion and linear Marangoni effect both result from entropic contributions of an underlying energy functional as shown for insoluble surfactants in Ref. [29].
The flow field generated by a non-uniform concentration field $c$ can be found by solving the Stokes equation in the fluid supplemented with the Marangoni stress boundary condition at the film surface $\mu\partial_{z}\mathbf{u}(x,y,z,t)\big{|}_{z=h}=-\frac{d\sigma}{dc}\mathbf{\nabla}c$, where $\sigma(c)$ is the concentration-dependent surface tension.
@@
This problem has been solved in Ref. [30] by assuming that the deformation of the film surface can be neglected, which is valid at small capillary numbers $\mu U/\sigma_{0}\ll 1$, where $\sigma_{0}$ is the surface tension of the bare liquid in the absence of surfactant. However, at large capillary number a complete description of the dynamics of a wet system would necessarily include the dynamics of film surface deformations. This will be addressed in a future study.
In particular, it was shown in Ref. [30] that for thick liquid layers of uniform thickness the surface velocity $\mathbf{U}$ induced by the Marangoni effect depends nonlocally on gradients of the surfactant density $c$ and can be represented as
$\mathbf{U}=-\frac{Eh}{4\pi^{2}c_{0}\mu}\int d^{2}k\frac{i\mathbf{k}\hat{c}(\sinh^{2}(|\mathbf{k}|h)-(|\mathbf{k}|h)^{2})}{|\mathbf{k}|h\sinh(2|\mathbf{k}|h)-2(|\mathbf{k}|h)^{2}}e^{i\mathbf{k}\cdot\mathbf{r}}.$ (2)
@@
Here, $E=-c_{0}(d\sigma(c)/dc)\big{|}_{c=c_{0}}>0$ is the Marangoni elasticity of the film surface that encodes the strength of change of the surface tension $\sigma(c)$ with $c$. Further, $\mathbf{k}=(k_{x},k_{y})$ is a 2D wave number and $\hat{c}(\mathbf{k},t)=4\pi^{2}\int d^{2}r\,c(\mathbf{r},t)e^{-i\mathbf{k}\cdot\mathbf{r}}$ is the Fourier transform of $c$.
@@
In long-wave approximation [17, 31, 32], valid when the thickness of the liquid film is small as compared to the characteristic horizontal length scale of the variation in surfactant density, i.e. $h\ll 2\pi/|\mathbf{k}|$, Eq. (2) reduces to the local relationship [17]
$\mathbf{U}(\mathbf{r},t)=-\frac{Eh}{4c_{0}\mu}\mathbf{\nabla}c(\mathbf{r},t).$ (3)
It is interesting to observe that the functional dependence of $\mathbf{U}$ on $c$ in equation (3) is identical to the one found for the velocity of a spherical solid colloidal particle transported at an interface by diffusiophoresis [26].
The uniform disordered state $\rho(\mathbf{r},\theta)=c_{0}/(2\pi)$ may become unstable to spatial and orientational perturbations, leading to a transition toward dynamic inhomogeneous ordered states including steady clusters or bands, and propagating waves as it was demonstrated previously for dry system [12, 13, 14]. One of the key observations derived from the numerical analysis of the Boltzmann equation (1) is that despite the short-range interactions due to collisions, which typically occur when the distance between the agents is comparable with their size, the global dynamics of the system exhibits a long-range order with a much larger characteristic scale. This allows us to replace the Boltzmann equations for the one-particle density $\rho(x,y,\theta,t)$ with a set of coupled hydrodynamic equations for certain coarse-grained order parameter fields that do not depend on the orientation angle $\theta$. As such, the Fourier transform of the density in the angular space is used
The orientation order on a large scale is well described by the first few order parameter fields
$\rho_{q}(\mathbf{r},t)=\int_{0}^{2\pi}e^{iq\theta}\rho(\mathbf{r},\theta,t)\,d\theta.$ (4)
In fact, for polar or polar-nematic alignment between the agents during collisions, the orientation order is well approximated by the first four dominant order parameters $\rho_{1,2,3,4}$ with all other order fields assumed to be negligibly small, i.e. $\rho_{q>4}=0$.
Our goal is to follow the closure scheme outlined in [13, 14] to derive a closed set of coarse-grained hydrodynamic equations for the density $c$ and the dominant order parameters $\rho_{1,2}$ for a wet system by including the advection velocity $\mathbf{U}$. Thus, we also assume the following scaling and truncation
$\begin{split}&|\mathbf{U}(c)|\sim v_{0},~~\partial_{t,x,y}\sim\epsilon,~~c-c_{0}\sim\epsilon,\\ &\rho_{1,2}\sim\epsilon,~~\rho_{3,4}\sim\epsilon^{2},~~\rho_{q>4}=0\end{split}$ (5)
@@
where $\epsilon\ll 1$ is the dimensionless super-criticality, i.e. the distance from the onset of the order-disorder transition. From the physical point of view, the above scaling implies a slow dynamics with a long-wave spatial orientation order dominated by polar $\rho_{1}$ and nematic $\rho_{2}$ order parameter fields. The scaling (5) automatically restricts the applicability of the hydrodynamic approximation to relatively small magnitudes and slow variations of $\rho_{1}$ and $\rho_{2}$. In consequence, similar to other weakly nonlinear approaches, the approximation is valid only in the vicinity of the order-disorder transition where the disordered state $c=c_{0}$ becomes unstable.
The closed form of the dynamic equations for the density $c$ and the complex-valued polar and nematic order parameter fields $\rho_{1}$ and $\rho_{2}$ is obtained in several steps as detailed in Appendix A. First, the Boltzmann equation is Fourier-transformed in the angular space and truncated at $q=4$, i.e. all order fields with $q>4$ are identically zero. Next, using scaling arguments (5), the fields $\rho_{3,4}$ are expressed in terms of $\rho_{1,2}$. Finally, after substituting $\rho_{3,4}$ into the equations for $c$ and $\rho_{1,2}$, the closed set of hydrodynamic equations is obtained as
$\partial_{t}c+\mathbf{\nabla}\cdot(\mathbf{U}c)$ $=$ $-\frac{v_{0}}{2}\left(\partial\rho_{1}^{*}+\partial^{*}\rho_{1}\right)+d\mathbf{\nabla}^{2}c$
$\partial_{t}\rho_{1}+\mathbf{\nabla}\cdot(\mathbf{U}\rho_{1})$ $=$ $-(\alpha_{0}+c\alpha_{1})\rho_{1}+\alpha_{2}\rho_{1}^{*}\rho_{2}-\alpha_{3}|\rho_{2}|^{2}\rho_{1}$ (6)
- $\frac{v_{0}}{2}(\partial c+\partial^{*}\rho_{2})+\gamma_{1}\rho_{2}^{*}\partial\rho_{2}+d\mathbf{\nabla}^{2}\rho_{1}$
$\partial_{t}\rho_{2}+\mathbf{\nabla}\cdot(\mathbf{U}\rho_{2})$ $=$ $(-\beta_{0}+c\beta_{1})\rho_{2}+\beta_{2}\rho_{1}^{2}-\beta_{3}|\rho_{2}|^{2}\rho_{2}$
@@
- $\gamma_{3}\rho_{1}^{*}\partial\rho_{2}-\gamma_{4}\partial^{*}(\rho_{1}\rho_{2})+d\mathbf{\nabla}^{2}\rho_{2}.$
Here, $\partial=\partial_{x}+i\partial_{y}$ and $\mathbf{\nabla}=(\partial_{x},\partial_{y})^{T}$ while the star indicates the complex conjugate. The system (6) is closed by Eq. (2) [or Eq. (3) in the long-wave limit] that relates $\mathbf{U}$ to the number density $c$. Note that we have retained translational diffusion with diffusivity $d$ to ensure numerical stability and consistency with the Marangoni term. For ease of comparison, system (6) is nondimensionalised using the scaling introduced in Ref. [14] for a dry system, i.e. when $\mathbf{U}=0$. Namely, time is scaled with the characteristic duration of the ballistic flight path of the agents $\tau=r^{-1}$, where $r$ is the rate of random orientation changes of the agents. The spatial coordinates are scaled with $v_{0}\tau$, and the one-particle density distribution $\rho$ is measured in units of $r/Dv_{0}$, where $D$ is the size of the agent particles. In dimensionless units, the dynamics of a wet system in the long-wave limit is described by Eqs. (6) with $v_{0}=1$ and $\mathbf{U}=-\text{Ma}\mathbf{\nabla}c$ where $\text{Ma}=Eh/(4v_{0}^{2}\mu\tau)$ represents the Marangoni parameter that controls the relative strength of the Marangoni flow as compared to the self-propulsion speed.
The coefficients $\alpha_{i}$ and $\beta_{i}$ in Eqs. (6) determine the properties of the spatially uniform states, while the $\gamma_{i}$ are responsible for the coupling between the gradients of polar and nematic order parameter fields.
As outlined in Appendix A, the closure relations for $\rho_{3,4}$ and, therefore, the coefficients in Eqs. (6) depend on the average density $c_{0}$, as well as on the microscopic parameters such as the statistics of the rotational diffusion, the effective interaction distance and the alignment rules associated with binary collisions. The behaviour of the system will therefore depend on changes in these microscopic parameters even if the average coarse-grained density is kept fixed. However, since our aim is to understand the effect of the self-induced Marangoni flow on the dynamics of the system on the coarse-grained level, we apply a semi-phenomenological approach as proposed in Ref. [14], i.e. we derive the governing equations by coarse graining but then pursue a parametric study on the continuum level. Accordingly, the obtained hydrodynamic equations (6) are studied for selected physically meaningful control parameters while all other parameters remain fixed. In particular, our main focus is the effect of the Marangoni number Ma and of the mean density $c_{0}$. Note that Ma can be varied independently at fixed density for a given set of the microscopic parameters.
Following Ref. [14], we choose the coefficients $\alpha_{i}$ and $\beta_{i}$ to correspond to a system that is known to feature several transitions in absence of Marangoni flow. When the mean number density $c_{0}$ is varied at fixed values of the coefficients $\alpha_{i}$, $\beta_{i}$ and $\gamma_{i}$, transitions occur between a uniform disordered state and a uniform ordered nematic state and as well between the latter and a uniform ordered polar-nematic state of mixed symmetry. The parameter values employed in the following are
@@
$\gamma_{1}$ $=$ $1,~\gamma_{2}=0.85,~\gamma_{3}=0.23,~\gamma_{4}=3.7.$ (7)
Coefficients $\alpha_{i}$, $\beta_{i}$ and $\gamma_{i}$ can be traced back to the microscopic parameters and the average density via the relations given in Appendix A. In fact, the values used here correspond to $c_{0}=0.08$ for the choice of the collision rules used in Ref. [14]. The translational diffusivity is varied between $d=10^{-2}$ and $d=10^{-1}$.